# On Inverse Problems, Parameter Estimation, and Domain Generalization

### v1

@@ 4.3.2 MMSE Estimator and posterior sampling
$\hat{\mathbf{x}}=E\{X|Y\}=\sum_{\mathbf{x}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ P(\mathbf{x}|\mathbf{y})=\sum_{\mathbf{x}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ \frac{P(\mathbf{y}|\mathbf{x})P(\mathbf{x})}{P(\mathbf{y})}$ (37)
@@
$\hat{\mathbf{x}}(\tilde{\mathbf{y}})=\mathbf{x}_{A}\frac{P_{X}(\mathbf{x}_{A})}{P_{Y}(\tilde{\mathbf{y}}=T(\mathbf{x}_{A}))}+\mathbf{x}_{B}\frac{P_{X}(\mathbf{x}_{B})}{P_{Y}(\tilde{\mathbf{y}}=T(\mathbf{x}_{B}))},$ (38)
@@
In other words, the reconstruction $\hat{\mathbf{x}}(\tilde{\mathbf{y}})$ is a weighted linear combination of the possible source signals explaining the given measurement. This outcome is also referred to as “regression to the mean” that often produces a significantly blurry outcome in image restoration tasks (Elad et al., 2023) (although not exactly representing the original term (Efron, 2011)). In our case study, obviously, the classification of the averaged $\hat{\mathbf{x}}(\tilde{\mathbf{y}})$ is under-determined. If by chance, for example, the averaged image belongs in one of the disjoint sets, then the image will be classified according to that class, regardless of the original image class. Recent work (Ohayon et al., 2024b) suggested to transport the outcome to the distribution of the source signal $p_{X}(\mathbf{x})$. However, in the case of restoration for the sake of downstream classification, this process essentially determines the estimated class. Let us denote,
$c_{A}=\frac{P(\theta_{1})P_{\theta_{1}}(\mathbf{x}_{A})}{P_{Y}(\tilde{\mathbf{y}})},\quad c_{B}=\frac{P(\theta_{2})P_{\theta_{2}}(\mathbf{x}_{B})}{P_{Y}(\tilde{\mathbf{y}})}.$ (39)
@@
$\hat{\mathbf{x}}=E\{X|Y\}=\sum_{\mathbf{x}\ :\ y=T(\mathbf{x})}\mathbf{x}\ P(\mathbf{x}|\mathbf{y})=\sum_{\mathbf{x}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ \frac{P(\mathbf{y}|\mathbf{x})P(\mathbf{x})}{P(\mathbf{y})}$ (40)
@@
$\hat{\mathbf{x}}(\tilde{\mathbf{y}})=\mathbf{x}_{A}\frac{P(\theta_{1})P(\mathbf{\tilde{y}}|\mathbf{x}_{A})P_{\theta_{1}}(\mathbf{x}_{A})}{P_{Y}(\tilde{\mathbf{y}}=T(\mathbf{x}_{A}))}+\mathbf{x}_{B}\frac{P(\theta_{2})P(\mathbf{\tilde{y}}|\mathbf{x}_{B})P_{\theta_{2}}(\mathbf{x}_{B}))}{P_{Y}(\tilde{\mathbf{y}}=T(\mathbf{x}_{B}))}.$ (41)
@@
$=\sum_{\mathbf{x}\in A_{X}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ \frac{P(\mathbf{y}|\mathbf{x})P(\mathbf{x})}{P(\mathbf{y})}+\sum_{\mathbf{x}\in B_{X}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ \frac{P(\mathbf{y}|\mathbf{x})P(\mathbf{x})}{P(\mathbf{y})}.$ (42)
@@
$p(\mathbf{x}|\mathbf{y})=\sum_{i}p(\mathbf{x}|\theta_{i},\mathbf{y})P(\theta_{i}|\mathbf{y})=\sum_{i}p_{\theta_{i}}(\mathbf{x}|\mathbf{y})P(\theta_{i}|\mathbf{y}).$ (43)
@@
$P(\theta_{i}|\mathbf{y})=q_{i}(\mathbf{y})=\frac{P(\theta_{i})p(\mathbf{y}|\theta_{i})}{p(\mathbf{y})},$ (44)
@@
4.3.3 Classification of a degraded/transformed signal
There are numerous examples of applications where given measurements are first processed to follow with a downstream task of classification, including: audio signals (Rosenbaum et al., 2025), speech and text (Dissen et al., 2024), communication (Erpek et al., 2020), computer vision (Pereg, 2025; Elad et al., 2023), biomedical imaging (Hanania et al., 2025; Pereg, 2023b), seismic imaging (Pereg et al., 2021, 2020b) and more. For example, in applications where the signal first goes through an inversion process, decompression, transformation into a different domain (such as: Radon transform) and so on. Then, the restored $\hat{\mathbf{x}}$ is classified to perform a diagnosis, or for other possible uses. Alternatively, some propose an end-to-end approach where the measurement is directly processed to perform the classification directly from the measurements. The discrete parameter $\theta$ represents the class to which the signal belongs as described above.
Note that in (30) we required that the inversion process $p(\hat{\mathbf{x}}|\mathbf{y})$ does not depend on $\theta$. This assumption is relevant since in many practical examples of real-world classification models that operate on a reconstructed image, the restoration model is agnostic to $\theta$, as demonstrated in Fig. 1. In other words, the restoration model is designed to yield an image that belongs to the general distribution of natural images (biomedical images, or any other specific distribution of signals). This is the case for biomedical signal classification after signal reconstruction/restoration. Namely, the inverse problem solved essentially implementing a process $p(\hat{\mathbf{x}}|\mathbf{y})$ (rather than $p_{\theta}(\hat{\mathbf{x}}|\mathbf{y})$), because the restoration model at that stage does not “know”’ the class to which the image belongs to. Theorem 6 is stating that although the inversion process may yield a visually enhanced image, the information in the measurement about the following classification can only be lost in the prior processing stage.
4.3.4 Segmentation of noisy (degraded) image
In many cases, instead of having one label for the entire image (or signal) we are interested in the label of each pixel (or point-value), a segment, or a patch. It is largely hypothesize and demonstrated in experimental results that denoising enhances segmentation results (Ye et al., 2023; Gruber et al., 2023; Zhao et al., 2023). Although, as we have shown above, in theory, due to the erroneous nature of the denoising process, equivalent accuracy could be obtained by performing segmentation directly from the noisy measurements.
5 Cramér-Rao Bound for Estimation After Data Processing
In a continuous setting, denote $\Delta_{Y}=J^{-1}_{Y}(\theta)$, and $\Delta_{\hat{X}}=J^{-1}_{\hat{X}}(\theta)$, such that by Cramér-Rao Bound
$\varepsilon_{Y}\triangleq E(\theta-\hat{\theta}_{Y})^{2}\geq\Delta_{Y},$ (46)
and
$\varepsilon_{\hat{X}}\triangleq E(\theta-\hat{\theta}_{\hat{X}})^{2}\geq\Delta_{\hat{X}},$ (47)
By definition of $\hat{X}=\mathcal{F}(Y)$, we have $I(\theta;Y)\geq I(\theta;\hat{X})$, and $J_{Y}(\theta)\geq J_{\hat{X}}(\theta)$ (Zamir, 1998). Thus, clearly, $\Delta_{Y}\leq\Delta_{\hat{X}}$. That is to say, the estimation error of a parameter $\theta$ of the observed measurement $Y$ is bounded by a smaller or equal quantity comparing with the estimation error of that parameter extracted by any deterministic representation that is obtained via processing of that measurement. Equality holds if $\hat{X}$ is a sufficient statistic relative to the family $\big{\{}f_{\theta}(y)\big{\}}$. Throughout the analysis we prove assuming $\theta$ is scalar. The generalization to a vector $\mathbf{\theta}$ is straightforward.
@@ 5 Cramér-Rao Bound for Estimation After Data Processing
Definition 8 (sufficient measurement)
@@
Definition 9 (sufficient representation)
@@
Theorem 10 (CRB for perfect perception estimator based on implicit parameter likelihood)
@@
Proof  Since $p_{\hat{X}|Y,\theta}(x|y,\theta)=p_{X|Y,\theta}(x|y,\theta)$, we have $p_{\hat{X}|\theta}(x|\theta)=p_{X|\theta}(x|\theta)$,
@@
[…]
Therefore $I(\theta,X)=I(\theta,\hat{X})$ and $J_{X}(\theta)=J_{\hat{X}}(\theta)$. From (12) and (13) it follows that $I(\theta;X)\geq I(\theta;Y)$ and $J_{X}(\theta)\geq J_{Y}(\theta)$ ((Zamir, 1998), Section I(6)). As stated above, due to DPI $H(\theta|\hat{X}=\mathcal{F}(Y))\geq H(\theta|Y)$, therefore, $I(\theta,Y)\geq I(\theta,\hat{X})$, which is only possible if $I(\theta;X)=I(\theta;Y)=I(\theta;\hat{X})$. Therefore, $\hat{X}$ is a sufficient statistic for $\theta$, and $J_{Y}(\theta)=J_{\hat{X}}(\theta)$. Equivalently, $\Delta_{Y}=\Delta_{\hat{X}}$.
Remark 6. Theorem 10 holds an implicit condition. As stated in the proof, by assumption it follows that $J_{X}(\theta)=J_{Y}(\theta)$, which is obeyed if $J_{X|Y}(\theta)=0$. In our Markov chain, $J_{Y|X}(\theta)=0$, because $p(y|x)$ does not depend on $\theta$. However, a perfect perception estimator $\hat{X}$ enforces a stricter requirement, that $J_{X}(\theta)=J_{Y}(\theta)=J_{\hat{X}}(\theta)$. In other words, for $\hat{X}$ to fully capture the information about the parameter $\theta$, $Y$ must be a sufficient representation of $X$. Mathematically speaking, it requires that $J_{X|Y}(\theta)=0$. This condition can be satisfied if $p_{X|Y}(x|y)$ does not depend on $\theta$. Alternatively, if $p_{X|Y,\theta}(x|y,\theta)$ is deterministic (Given $Y$, $X$ is deterministic), .i.e. $p_{X|Y,\theta}(x|y,\theta)$ is invertible, the FI is perfectly preserved in $Y$, and hence also in $\hat{X}$.
@@
Theorem 11 (CRB for perfect perception estimator based on common parameter prior)
@@
Proof  Since $p_{\hat{X}|Y,\theta}(x|y,\theta=\theta_{k})=p_{X|Y,\theta}(x|y,\theta=\theta_{k})$ for a fixed (unknown) $\theta=\theta_{k}$, we have $p_{X|\theta}(x|\theta=\theta_{k})=p_{\hat{X}|\theta}(x|\theta=\theta_{k})$, since
@@
Proposition 12 (CRB for perfect prior perception estimator)
Proof  $p_{X}(x)=p_{\hat{X}}(x)$ does not necessarily guarantees that $I(\theta,X)=I(\theta,\hat{X})$ nor that $J_{X}(\theta)=J_{\hat{X}}(\theta)$. From (12) and (13) it follows that $I(\theta;X)\geq I(\theta;Y)$. As stated above, due to DPI $H(\theta|\hat{X}=\mathcal{F}(Y))\geq H(\theta|Y)$, therefore, $I(\theta,Y)\geq I(\theta,\hat{X})$.
@@
Remark 10. Theorems 10-11 imply that the recovery system or method knows or is able to implicitly detect $\theta$, which is in practice, very challenging to accomplish.
@@
An immediate consequence of the DPI is that information about $\theta$ that is lost in $Y$ cannot be recovered in $\hat{X}$, as $\hat{x}=\mathcal{F}(\hat{y})$ is the outcome of the processing of $Y$. Namely, it holds that
$I(\theta;X)\geq I(\theta;Y)\geq I(\theta;\hat{X}),$ (48)
with equality if and only if $\hat{X}$ is a sufficient statistic of $Y$, and $Y$ is a sufficient representation of $X$. Consequently, the second case we consider is $I(\theta;X)\geq I(\theta;Y)=I(\theta;\hat{X})$. In this case, $\hat{X}=\mathcal{F}(\hat{Y})$ is a sufficient statistic of $Y$, $X-\hat{X}-Y$ and $\Delta_{Y}=\Delta_{\hat{X}}$. By Neyman’s factorization criterion for a sufficient statistic it holds that $P_{\theta}(Y)=h(Y)Q_{\theta}(\hat{X})$. Note that in this case, regardless of how close $P_{\hat{X}}$ is to $P_{X}$, the estimation error using the measurement directly or its processing is determined by the distribution of the sufficient statistic relative to $Y$. Therefore, when $\hat{X}$ is not a sufficient statistic of $Y$, we have our third case, namely $\Delta_{Y}>\Delta_{\hat{X}}$, since $J_{Y}(\theta)>J_{\hat{X}}(\theta)$. Table 1 summarizes the above analysis into 3 cases.
@@
$p_{\hat{X}|Y,\theta}(x|y,\theta)=p_{X|Y,\theta}(x|y,\theta)$ $p_{\hat{X}|\theta}(x|\theta)=p_{X|\theta}(x|\theta)$ $I(\theta;X)=I(\theta;Y)=I(\theta;\hat{X})$ $\Delta_{Y}=\Delta_{\hat{X}}$
$p_{\hat{X}|Y,\theta}(x|y,\theta)\neq p_{X|Y,\theta}(x|y,\theta)$ $p_{\hat{X}|\theta}(x|\theta)\neq p_{X|\theta}(x|\theta)$ $I(\theta;X)\geq I(\theta;Y)=I(\theta;\hat{X})$ $\Delta_{Y}\leq\Delta_{\hat{X}}$
$p_{\hat{X}|Y,\theta}(x|y,\theta)\neq p_{X|Y,\theta}(x|y,\theta)$ $p_{\hat{X}|\theta}(x|\theta)\neq p_{X|\theta}(x|\theta)$ $I(\theta;X)\geq I(\theta;Y)>I(\theta;\hat{X})$ $\Delta_{Y}<\Delta_{\hat{X}}$
@@
As mentioned above, it has been shown that the minimal deviation between $p_{X}$ and $p_{\hat{X}}$ that can be obtained by an estimator increases as the mean distortion decreases. In other words, restoration algorithms attempting to minimize a distortion measure inherently contradict an attempt to improve perceptual quality. In the context of this work, in accordance with Theorems 10-11, the accuracy of the parameter estimation is determined by the accuracy in retrieving the conditional source’s distribution with respect to the parameter estimated, rather than successfully minimizing a distortion measure. In other words, an estimator $\hat{X}$ of $X$ that is designed to minimize a distortion measure rather than find the closest source distribution, is bound to be erroneous in terms of the following parameter estimation. Note that by Theorems 10-11, an estimator that produces the “wrong” $\hat{X}$ in terms of the quality measure between $X=x$ and $\hat{X}=\hat{x}$, yet the estimation still belongs to the source conditional distribution $\hat{X}|\theta\in p_{X|\theta}(\cdot)$, would be as reliable as the “correct” reconstruction. For example, suppose we would like to estimate the degree of elasticity of the retina from and OCT speckled image. As a first step, we despeckle the image, and our model accidentally produces a despeckled image which does not correspond with the input, yet is an OCT image of a possible retina of a patient corresponding to the same elasticity measure. Therefore the erroneous nature of the reconstruction system w.r.t the quality measure between $X=x$ and $\hat{X}=\hat{x}$ is irrelevant to the following downstream task of estimating $\theta$. That said, as pointed out in Theorem 12, an estimator that attempts to recover a signal that belongs to the general distribution of the ground-truth $p_{X}\forall\theta$ (e.g., (Ohayon et al., 2024b)) is not guaranteed to produce a better estimation of the desired parameter from which that ground truth belongs to. Overall, even if the recovered signal belongs to the true conditional ground-truth distribution the accuracy would still be as good as the accuracy obtained by estimating $\theta$ directly from the measurement.

### v2

@@ 4.3.2 MMSE Estimator and posterior sampling
(41) $\hat{\mathbf{x}}=E\{X|Y\}=\sum_{\mathbf{x}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ P(\mathbf{x}|\mathbf{y})=\sum_{\mathbf{x}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ \frac{P(\mathbf{y}|\mathbf{x})P(\mathbf{x})}{P(\mathbf{y})}$
@@
(42) $\hat{\mathbf{x}}(\tilde{\mathbf{y}})=\mathbf{x}_{A}\frac{P_{X}(\mathbf{x}_{A})}{P_{Y}(\tilde{\mathbf{y}}=T(\mathbf{x}_{A}))}+\mathbf{x}_{B}\frac{P_{X}(\mathbf{x}_{B})}{P_{Y}(\tilde{\mathbf{y}}=T(\mathbf{x}_{B}))},$
@@
[…]
(43) $c_{A}=\frac{P(\theta_{1})P_{\theta_{1}}(\mathbf{x}_{A})}{P_{Y}(\tilde{\mathbf{y}})},\quad c_{B}=\frac{P(\theta_{2})P_{\theta_{2}}(\mathbf{x}_{B})}{P_{Y}(\tilde{\mathbf{y}})}$.
@@
(44) $\hat{\mathbf{x}}=E\{X|Y\}=\sum_{\mathbf{x}\ :\ y=T(\mathbf{x})}\mathbf{x}\ P(\mathbf{x}|\mathbf{y})=\sum_{\mathbf{x}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ \frac{P(\mathbf{y}|\mathbf{x})P(\mathbf{x})}{P(\mathbf{y})}$
@@
(45) $\hat{\mathbf{x}}(\tilde{\mathbf{y}})=\mathbf{x}_{A}\frac{P(\theta_{1})P(\mathbf{\tilde{y}}|\mathbf{x}_{A})P_{\theta_{1}}(\mathbf{x}_{A})}{P_{Y}(\tilde{\mathbf{y}}=T(\mathbf{x}_{A}))}+\mathbf{x}_{B}\frac{P(\theta_{2})P(\mathbf{\tilde{y}}|\mathbf{x}_{B})P_{\theta_{2}}(\mathbf{x}_{B}))}{P_{Y}(\tilde{\mathbf{y}}=T(\mathbf{x}_{B}))}.$
@@
(46) $=\sum_{\mathbf{x}\in A_{X}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ \frac{P(\mathbf{y}|\mathbf{x})P(\mathbf{x})}{P(\mathbf{y})}+\sum_{\mathbf{x}\in B_{X}\ :\ \mathbf{y}=T(\mathbf{x})}\mathbf{x}\ \frac{P(\mathbf{y}|\mathbf{x})P(\mathbf{x})}{P(\mathbf{y})}.$
@@
(47) $P(\mathbf{x}|\mathbf{y})=\sum_{i}P(\mathbf{x}|\theta_{i},\mathbf{y})P(\theta_{i}|\mathbf{y})=\sum_{i}P_{\theta_{i}}(\mathbf{x}|\mathbf{y})P(\theta_{i}|\mathbf{y}).$
@@
(48) $P(\theta_{i}|\mathbf{y})=q_{i}(\mathbf{y})=\frac{P(\theta_{i})P(\mathbf{y}|\theta_{i})}{P(\mathbf{y})},$
@@
(49) $p(\mathbf{x}|\mathbf{y})=\int p(\mathbf{x}|\theta,\mathbf{y})p(\theta|\mathbf{y})d\theta=\int p_{\theta}(\mathbf{x}|\mathbf{y})p(\theta|\mathbf{y})d\theta.$
4.3.3 Classification of a degraded/transformed signal
There are numerous examples of applications where given measurements are first processed to follow with a downstream task of classification, including: audio signals [Rosenbaum:2025], speech and text [Dissen:2024], communication [Clancy:2020], computer vision [Pereg:2024B, Elad:2023], biomedical imaging [Hanania:2025, Pereg:2023B], seismic imaging [Pereg:2021, Pereg:2020b] and more. For example, in applications where the signal first goes through an inversion process, decompression, transformation into a different domain (such as: Radon transform) and so on. Then, the restored $\hat{\mathbf{x}}$ is classified to perform a diagnosis, or for other possible uses. Alternatively, some propose an end-to-end approach where the measurement is directly processed to perform the classification directly from the measurements. The discrete parameter $\theta$ represents the class to which the signal belongs as described above.
Note that in (34) we required that the inversion process $p(\hat{\mathbf{x}}|\mathbf{y})$ does not depend on $\theta$. This assumption is relevant since in many practical examples of real-world classification models that operate on a reconstructed image, the restoration model is agnostic to $\theta$, as demonstrated in Fig. 1. In other words, the restoration model is designed to yield an image that belongs to the general distribution of natural images (biomedical images, or any other specific distribution of signals). This is the case for biomedical signal classification after signal reconstruction/restoration. Namely, the solved inverse problem is essentially implementing a process $p(\hat{\mathbf{x}}|\mathbf{y})$ (rather than $p_{\theta}(\hat{\mathbf{x}}|\mathbf{y})$), because the restoration model at that stage does not “know”’ the class to which the image belongs to. Theorem 4.3 is stating that although the inversion process may yield a visually enhanced image, the information in the measurement about the following classification can only be lost in the prior processing stage.
4.3.4 Segmentation of noisy (degraded) image
In many cases, instead of having one label for the entire image (or signal) we are interested in the label of each pixel (or point-value), a segment, or a patch. It is largely hypothesize and demonstrated in experimental results that denoising enhances segmentation results [Ye:2023b, Gruber:2023, Zhao:2023]. Although, as we have shown above, in theory, due to the erroneous nature of the denoising process, equivalent accuracy could be obtained by performing segmentation directly from the noisy measurements.
5 Cramér-Rao Bound for Estimation After Data Processing
In a continuous setting, denote $\Delta_{Y}=J^{-1}_{Y}(\theta)$, and $\Delta_{\hat{X}}=J^{-1}_{\hat{X}}(\theta)$, such that by Cramér-Rao Bound, for unbiased estimators we have
(50) $\varepsilon_{Y}\triangleq E(\theta-\hat{\theta}_{Y})^{2}\geq\Delta_{Y},$
and
(51) $\varepsilon_{\hat{X}}\triangleq E(\theta-\hat{\theta}_{\hat{X}})^{2}\geq\Delta_{\hat{X}}$,
@@ 5 Cramér-Rao Bound for Estimation After Data Processing
Definition 5.1 (sufficient measurement).
@@
Definition 5.2 (sufficient representation).
@@
Theorem 5.3 (CRB for perfect perception estimator based on implicit parameter likelihood).
@@
Proof 5.4.
Since $p_{\hat{X}|Y,\theta}(x|y,\theta)=p_{X|Y,\theta}(x|y,\theta)$, we have $p_{\hat{X}|\theta}(x|\theta)=p_{X|\theta}(x|\theta)$,
$p_{\hat{X}|\theta}(x|\theta)=\int p_{\hat{X}|Y,\theta}(x|y,\theta)p_{Y|\theta}(y|\theta)dy=\int p_{X|Y,\theta}(x|y,\theta)p_{Y|\theta}(y|\theta)dy=p_{X|\theta}(x|\theta).$
$p_{\hat{X}}(x)=\int p_{\hat{X}|\theta}(x|\theta)p_{\Theta}(\theta)d\theta=\int p_{X|\theta}(x|\theta)p_{\Theta}(\theta)d\theta=p_{X}(x).$
@@
Remark 6. Theorem 5.3 holds an implicit condition. As stated in the proof, by assumption it follows that $J_{X}(\theta)=J_{Y}(\theta)$, which is obeyed if $J_{X|Y}(\theta)=0$. In our Markov chain, $J_{Y|X}(\theta)=0$, because $p(y|x)$ does not depend on $\theta$. However, a perfect perception estimator $\hat{X}$ enforces a stricter requirement, that $J_{X}(\theta)=J_{Y}(\theta)=J_{\hat{X}}(\theta)$. In other words, for $\hat{X}$ to fully capture the information about the parameter $\theta$, $Y$ must be a sufficient representation of $X$. Mathematically speaking, it requires that $J_{X|Y}(\theta)=0$. This condition can be satisfied if $p_{X|Y}(x|y)$ does not depend on $\theta$.Alternatively, if $p_{X|Y,\theta}(x|y,\theta)$ is deterministic (Given $Y$, $X$ is deterministic), .i.e. $p_{X|Y,\theta}(x|y,\theta)$ is invertible, the FI is perfectly preserved in $Y$, and hence also in $\hat{X}$.
A simple example for an invertible non-deterministic mapping is a noisy constellation $\mathbf{y}=\mathbf{x}+\mathbf{e}$, where $\mathcal{X}^{n}=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{M}\}$ and $\mathcal{E}^{n}=\{\mathbf{e}_{1},\mathbf{e}_{2},...,\mathbf{e}_{P}\}$ are independent and both uniformly distributed, where we further assume that the map $\mathbf{y}\rightarrow\mathbf{x}$ is surjective. In other words, the noise is stochastic yet the measurements could still be reliably mapped back to their origin.
@@
Theorem 5.5 (CRB for perfect perception estimator based on common parameter prior).
@@
Proof 5.6.
Since $p_{\hat{X}|Y,\theta}(x|y,\theta=\theta_{k})=p_{X|Y,\theta}(x|y,\theta=\theta_{k})$ for a fixed (unknown) $\theta=\theta_{k}$, we have $p_{X|\theta}(x|\theta=\theta_{k})=p_{\hat{X}|\theta}(x|\theta=\theta_{k})$, since
$p_{\hat{X}|\theta}(x|\theta=\theta_{k})=\int p_{\hat{X}|Y,\theta}(x|y,\theta=\theta_{k})dy=\int p_{X|Y,\theta}(x|y,\theta=\theta_{k})dy=p_{X|\theta}(x|\theta=\theta_{k}).$
Proposition 5.7 (CRB for perfect prior perception estimator).
Proof 5.8.
@@
Remark 10. Theorems 5.3-5.5 imply that the recovery system or method knows or is able to implicitly detect $\theta$, which is in practice, very challenging to accomplish.
@@
[…]
@@
[…]
@@
As mentioned above, it has been shown that the minimal deviation between $p_{X}$ and $p_{\hat{X}}$ that can be obtained by an estimator increases as the mean distortion decreases. In other words, restoration algorithms attempting to minimize a distortion measure inherently contradict an attempt to improve perceptual quality. In the context of this work, in accordance with Theorems 5.3-5.5, the accuracy of the parameter estimation is determined by the accuracy in retrieving the conditional source’s distribution with respect to the parameter estimated, rather than successfully minimizing a distortion measure. In other words, an estimator $\hat{X}$ of $X$ that is designed to minimize a distortion measure rather than find the closest source distribution, is bound to be erroneous in terms of the following parameter estimation. Note that by Theorems 5.3-5.5, an estimator that produces the “wrong” $\hat{X}$ in terms of the quality measure between $X=x$ and $\hat{X}=\hat{x}$, yet the estimation still belongs to the source conditional distribution $\hat{X}|\theta\in p_{X|\theta}(\cdot)$, would be as reliable as the “correct” reconstruction. For example, suppose we would like to estimate the degree of elasticity of the retina from an OCT speckled image. As a first step, we despeckle the image, and our model accidentally produces a despeckled image which does not correspond with the input, yet is an OCT image of a possible retina of a patient corresponding to the same elasticity measure. Therefore the erroneous nature of the reconstruction system w.r.t the quality measure between $X=x$ and $\hat{X}=\hat{x}$ is irrelevant to the following downstream task of estimating $\theta$. That said, as pointed out in Theorem 5.7, an estimator that attempts to recover a signal that belongs to the general distribution of the ground-truth $p_{X}\forall\theta$ (e.g., [Ohayon:2024a]) is not guaranteed to produce a better estimation of the desired parameter from which that ground truth belongs to. Overall, even if the recovered signal belongs to the true conditional ground-truth distribution the accuracy would still be as good as the accuracy obtained by estimating $\theta$ directly from the measurement.